OpenStudy (anonymous):

A hot wire radiates heat at 100 Watts. If its temperature measured in degrees Kelvin is doubled, then the power radiated will be what? 800Watts or 1600Watts

OpenStudy (anonymous):

Power from a radiative body is a function of temperature to the fourth power.

OpenStudy (jamesj):

Exactly. By the Stefan-Boltzman law, \[ P_0 = A \epsilon \sigma T^4 \] Hence if you double T, \[ P = A \epsilon \sigma (2T)^4 = 2^4 (A \epsilon \sigma T^4) = 16 P_0 \]